# Convergence Rates for the Maximum A Posteriori Estimator in PDE-Regression Models with Random Design

### v1

@@ 1. C7)
    Inverse continuity modulus: Additionally to 2.3, define for any $\delta\in\mathbb{R}_{>0}$ the set
    $\Lambda_{\delta}:=\left\{(\theta_{1},\theta_{2})\in\tilde{\Theta}^{2}:\,\left\|\theta_{1}\right\|_{H^{\alpha}(\mathcal{O},W)}+\left\|\theta_{2}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{M},\,\left\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq\delta\right\}$
@@
    $\sup\left\{\left\|\theta_{1}-\theta_{2}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}:\,(\theta,\theta^{\prime})\in\Lambda_{\delta}\right\}\leq C_{\operatorname{M}}\cdot\delta^{\tau}.$
@@
Under 2.3 and 2.9 and under the conditions formulated in Theorem 2.6, fory every $\bar{c}\in\mathbb{R}_{>0}$, there exists $c_{7}=c_{7}(C_{\operatorname{M}},\bar{c},\left\|\theta_{\mathfrak{o}}\right\|_{H^{\alpha}(\mathcal{O},W)},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\left\{\left\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\right\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c_{7}\cdot\mathfrak{r}_{N}^{\tau},\,\left\|\hat{\theta}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{7}\right\}\right)\geq 1-c_{8}\cdot\exp(-\bar{c}N\mathfrak{r}_{N}^{2})$
@@
3   3Application: PDE constrained regression models
We apply the results from Section 2 on the inverse problems described by the PDE Eq. 1.2 and derive corresponding convergence rates for the parameter $f$. We further discuss the applicability of the setup for other PDE models. Throughout this section, we are interested in estimating the unknown scalar function $f:\mathcal{O}\to\mathbb{R}$ and corresponding solutions $u=u_{f}:\mathcal{O}\to\mathbb{R}$ of the Darcy problem. For any $d\in\mathbb{N}$, integer $\alpha>\frac{d}{2}$, a fixed constant $f_{\min}\in\mathbb{R}_{>0}$, and denoting by $\partial_{n}$ the outward pointing normal derivative, define the parameter space
$\mathcal{F}:=\mathcal{F}_{\alpha,f_{\min}}:=\left\{f\in H^{\alpha}(\mathcal{O}):\,f>f_{\min}\text{ on $\mathcal{O}$},f|_{\partial\mathcal{O}}\equiv 1,\partial^{j}_{n}f\equiv 0,j\in\left\llbracket\alpha-1\right\rrbracket\right\},$ (3.1)
and its subclass $\mathcal{F}(\operatorname{L}):=\mathcal{F}\cap H^{\alpha}(\mathcal{O},\operatorname{L})$, $\operatorname{L}\in\mathbb{R}_{>0}$. Let $g\in C^{\infty}(\mathcal{O})$ be a known heat source function. If $f\in\mathcal{F}_{\alpha,f_{\min}}$ for some $\alpha>\frac{d}{2}+1$ and $f_{\min}\in\mathbb{R}_{>0}$, then well-known PDE theory (see Gilbarg and Trudinger [2001] for instance) provides the existence and uniquess of a classical solution $u\in C^{0}(\overline{\mathcal{O}})\cap C^{2+\nu}(\mathcal{O})$ of the Dirichlet boundary value problem
$\begin{dcases}\operatorname{div}(f(x)\nabla u(x))&=g(x)\text{ on $\mathcal{O}$,}\\ \hfill u(x)&=0\text{ on $\partial\mathcal{O}$,}\end{dcases}$ (3.2)
for some (small) $\nu\in\mathbb{R}_{>0}$. In other words, the solution operator $\mathcal{F}\ni f\mapsto G^{D}(f)=:u_{f}\in\mathbb{L}^{2}(\mathcal{O})$ is always well-defined. The nonlinearity of $G^{D}$ becomes more clear when considering the Feynmann-Kac formula, which reads for each $x\in\mathcal{O}$ as
$u_{f}(x)=-\frac{1}{2}\mathbb{E}^{x}\left[\int_{0}^{\tau_{\mathcal{O}}}g(X_{s}^{f})\mathrm{d}s\right],$
where $(X_{t}^{f})_{t\in\mathbb{R}_{\geq 0}}$ is a Markov process starting at $x$ with associated exit time $\tau_{\mathcal{O}}$. We refer to Bass [1998] for an extensive overview and more details. Following Eq. 2.1, we then observe
$Y_{i}=G^{D}(f_{\mathfrak{o}})(X_{i})+\sigma\varepsilon_{i},i\in\left\llbracket N\right\rrbracket,$ (3.3)
@@
    $\sup_{x\in\mathbb{R}}\left\{\left|\Psi^{(k)}(x)\right|\right\}<\infty,$
@@
For examples and typical choices of (regular) link functions in practise, we refer to Nickl et al. [2020] and Nickl and Wang [2024]. In the notation of Section 2, throughout this section we set $\mathcal{Z}=\mathcal{O}$, $\mathscr{Z}:=\mathscr{B}_{\mathcal{O}}$ and $\zeta=\mathcal{L}_{\mathcal{O}}$ (recalling that $\operatorname{vol}(\mathcal{O})=1$!). We set $\tilde{\Theta}:=\Theta:=\left\{\Psi^{-1}\circ f:\,f\in\mathcal{F}\right\}=H^{\alpha}_{c}(\mathcal{O})$ (the pulled-back) parameter space, and
@@
$J_{\mathfrak{r},N}:\mathcal{F}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\,f\mapsto-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-G^{D}(f)(X_{i})|^{2}-\frac{\mathfrak{r}^{2}}{2}\left\|\Psi^{-1}\circ f\right\|_{H^{\alpha}(\mathcal{O})}^{2}$ (3.4)
for any $\mathfrak{r}\in\mathbb{R}_{>0}$. In the notation of Eq. 2.2, we have $J_{\mathfrak{r},N}[f]=\mathscr{J}_{\mathfrak{r},N}[\theta]$ for any $f=f_{\theta}\in\mathcal{F}$, such that maximizing $J_{\mathfrak{r},N}$ over $\mathcal{F}$ is equivalent to maximizing $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta=H^{\alpha}_{c}(\mathcal{O})$. Any pairs of maximizer will be denoted by
@@
$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{1},f_{2}):=\left\|G(f_{1})-G(f_{2})\right\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}^{2}+\mathfrak{r}^{2}\left\|\Psi^{-1}\circ f_{1}\right\|_{H^{\alpha}(\mathcal{O})}^{2}.$
@@
Let $\mathcal{F}$ be as in Eq. 3.1 for some $\alpha>1+\frac{5}{2}d$ and $f_{\min}\in\mathbb{R}_{>0}$. Let $G^{D}(f)=u_{f}$ denote the unique solution of Eq. 3.2 and let $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$ from Eq. 3.3 for some $f_{\mathfrak{o}}\in\mathcal{F}$. Let $\Psi$ be a regular link function. Let $J_{\mathfrak{r}_{N},N}$ be as in Eq. 3.4 with
@@
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{P}_{f}^{N}\left(\left\{\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f)\geq\operatorname{M}\mathfrak{r}_{N}^{2}\right\}\right)\right\}\leq c_{2}\exp\left(-c_{3}\operatorname{M}^{2}N\mathfrak{r}_{N}^{2}\right).$ (3.5)
@@
    Assume additionally that $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\bar{c}\in\mathbb{R}_{>0}$ exists $c_{2}=c_{2}(\alpha,d,\mathcal{O},f_{\min},g_{\min},f_{\mathfrak{o}},\sigma)$ large enough and $c_{3}\in\mathbb{R}_{>0}$, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\left\{\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f_{\mathfrak{o}})\right\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{2}\mathfrak{r}_{N}^{\frac{\alpha-1}{\alpha+1}},\,\left\|\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\right\|_{H^{\alpha}(\mathcal{O})}\leq c_{2}\right\}\right)\geq 1-c_{3}\exp(-c_{1}N\mathfrak{r}_{N}^{2}).$
@@
Grant the conditions formulated in Theorem 3.2.
@@
    (Prediction Error) For every $\operatorname{L}\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$, there exists a finite constant $c_{4}=c_{4}(\alpha,\beta,d,\mathcal{O},f_{\min},L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f)\right\|_{H^{\beta}(\mathcal{O})}\right]\right\}\leq c_{4}\cdot N^{-\frac{\alpha+1-\beta}{2(\alpha+1)+d}}.$
@@
    (Estimation Error) Let $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\operatorname{L}\in\mathbb{R}_{>0}$, there exists $c_{5}=c_{5}(g_{\min},f_{\min},\mathcal{O},d,\alpha,L,\left\|g\right\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\left\|\hat{f}_{\textsc{MAP}}-f\right\|_{\mathbb{L}^{2}(\mathcal{O})}\right]\right\}\leq c_{5}\cdot N^{-\frac{\alpha-1}{2(\alpha+1)+d}}.$
Remark § 3.4 (Rate-optimality).
The convergence rates for the prediction risk presented in Theorem 3.3 correspond to the usual interpretation of statistical inverse problems: the smoothness of the parameter of interest $f$ is represented by the parameter $\alpha$. The ill-posedness of the underlying problem by $\kappa=1$, which corresponds to the smoothing property of $G^{D}$ (Indeed, Lemma 23 in Nickl et al. [2020] shows that if $f$ is $\alpha$-smooth, then $G^{D}(f)$ is $\alpha+1$-smooth). The prediction risk is considered with respect to the $H^{\beta}$-norm (which can be interpreted as differentiating $\beta$ times), which yields the characteristic numerator $\alpha+1-\beta$ of the rate. The rates of convergence here coincide with the rates obtained in Nickl et al. [2020] for the Darcy problem. There it is also shown that the rate of convergence of the prediction risk is minimax-optimal for the choice $\beta=2$, which also holds true in the random design model, as the proof can be easily adopted.
The proofs of Theorem 3.2 and Theorem 3.3 are given in detail in Appendix A. They are based on verifying that the forward map $\mathscr{G}^{D}$ satisfies C1)-C4) formulated in 2.3 and 2.9 for suitable parameters $\alpha,\kappa,\gamma_{1},\gamma_{2},\eta_{1}$ and $\eta_{2}$, with $V=W=\mathbb{R}$, and an application of the results in Section 2.
@@
The general framework presented in Section 2 does not only apply to the Darcy problem, it promises to consider various PDE-related problems. As already brielfly discussed in Remark 2.4, the results in Bohr and Nickl [2023] concerning the MAP estimate are special cases by setting $\eta_{1}=\eta_{2}=\gamma_{2}=0$. Hence, the solution map of the time-independent Schrödinger equation satisfies also meets the requirements of 2.3 and 2.9 taking the results of Nickl et al. [2020] also into account. Moreover, very analogous and elementary computations show that the solution map of the heat equation studied in Kekkonen [2022] also satisfies the conditions in 2.3 and 2.9 with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
In very recent work, Nickl and Titi [2024] are considering a non-linear Bayesian data assimilation model for the periodic $d=2$-dimensional Navier Stokes equation with unknown initial condition. In this setting $u=u_{\theta}:\mathcal{O}\times(0,T)\to\mathbb{R}^{2}$ solves the non-linear evolution equation
@@
where $\theta$ is the initial condition of the system, $f$ a forcing term and $\nu\in\mathbb{R}_{>0}$ a viscosity parameter. For this type of equation in two dimensions a complete solution theory already exists. In particular, if the initial condition $\theta$ is contained in some specific function class $\tilde{\mathcal{F}}$, the solution operator ~$\tilde{\mathcal{F}}\ni\theta\mapsto G^{NS}(\theta)=u_{f}\in\mathbb{L}^{2}(\mathcal{O}\times(0,T))$ is always well-defined. Based on a data set $D_{N}\in(\mathbb{R}^{2}\times\mathcal{O}\times(0,T))^{N}$ arising from Eq. 2.1, it seems feasible to apply the theory from Section 2. Indeed, using the theory presented in Nickl and Titi [2024], elementary computations show that the solution map is $(\alpha,0,0)$-regular in $\mathbb{L}^{2}$-norm and uniformly bounded, thus satisfying C1) and C2) from 2.3. C7) from 2.9 can also be shown with $\tau=1$. If C3) of 2.3 is additionally satisfied, one would obtain rates of the form $N^{-\frac{\alpha}{\alpha+2}}$ for estimating $\theta$ both in prediction and estimation error.
Appendix A Appendix
   Proofs of Section 2

### v3

@@ 1. C7)
    Inverse continuity modulus: Additionally to 2.3, define for any $\delta\in\mathbb{R}_{>0}$
    $\Lambda_{\delta}:=\left\{(\theta_{1},\theta_{2})\in\tilde{\Theta}^{2}:\,\hskip-2.84526pt\|\theta_{1}\|_{H^{\alpha}(\mathcal{O},W)}+\|\theta_{2}\|_{H^{\alpha}(\mathcal{O},W)}\leq\operatorname{M},\,\|\mathscr{G}(\theta_{1})-\mathscr{G}(\theta_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z},V)}\leq\delta\right\}$
@@
    $\sup\left\{\|\theta_{1}-\theta_{2}\|_{\mathbb{L}^{2}(\mathcal{O},W)}:\,(\theta,\theta^{\prime})\in\Lambda_{\delta}\right\}\leq C_{\operatorname{M}}\cdot\delta^{\tau}.$
@@
Under 2.3 and 2.9 and under the conditions formulated in theorem 2.6, fory every $\bar{c}\in\mathbb{R}_{>0}$, there exists $c_{7}(C_{\operatorname{M}},\bar{c},\|\theta_{\mathfrak{o}}\|_{H^{\alpha}(\mathcal{O},W)},\texttt{P})\in\mathbb{R}_{>0}$ large enough, such that any maximizer $\hat{\theta}_{\textsc{MAP}}$ of $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta$ satisfies
$\mathbb{P}_{\theta_{\mathfrak{o}}}^{N}\left(\|\hat{\theta}_{\textsc{MAP}}-\theta_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O},W)}\leq c_{7}\cdot\mathfrak{r}_{N}^{\tau},\,\|\hat{\theta}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O},W)}\leq c_{7}\right)\geq 1-c_{8}\cdot\exp(-\bar{c}\cdot N\mathfrak{r}_{N}^{2})$
@@
3 Application: PDE constrained regression models
We apply the results from section 2 on the inverse problems described by the PDE eq. 2 and derive corresponding convergence rates for the parameter $f$. We further discuss the applicability of the setup for other PDE models. Throughout this section, we are interested in estimating the unknown scalar function $f:\mathcal{O}\to\mathbb{R}$ and corresponding solutions $u=u_{f}:\mathcal{O}\to\mathbb{R}$ of the Darcy problem. For any $d\in\mathbb{N}$, integer $\alpha>\frac{d}{2}$, a fixed constant $f_{\min}\in\mathbb{R}_{>0}$, and denoting by $\partial_{n}$ the outward pointing normal derivative, define the parameter space
$\mathcal{F}:=\mathcal{F}_{\alpha,f_{\min}}:=\left\{f\in H^{\alpha}(\mathcal{O}):\,f>f_{\min}\text{ on $\mathcal{O}$},f|_{\partial\mathcal{O}}\equiv 1,\partial^{j}_{n}f\equiv 0,j\in\left\llbracket\alpha-1\right\rrbracket\right\},$ (12)
and its subclass $\mathcal{F}(\operatorname{L}):=\mathcal{F}\cap H^{\alpha}(\mathcal{O},\operatorname{L})$, $\operatorname{L}\in\mathbb{R}_{>0}$. Let $g\in C^{\infty}(\mathcal{O})$ be a known heat source function. If $f\in\mathcal{F}_{\alpha,f_{\min}}$ for some $\alpha>1+\frac{d}{2}$ and $f_{\min}\in\mathbb{R}_{>0}$, then well-known PDE theory (see Gilbarg and Trudinger (2001) for instance) provides the existence and uniquess of a classical solution $u\in C^{0}(\overline{\mathcal{O}})\cap C^{2+\nu}(\mathcal{O})$ of the Dirichlet boundary value problem
$\begin{dcases}\hfill\operatorname{div}(f(x)\cdot\nabla u(x))&=g,\text{ on $\mathcal{O}$},\\ \hfill u(x)\hskip 2.84526pt&=0,\text{ on $\partial\mathcal{O}$},\end{dcases}$ (13)
for some (small) $\nu\in\mathbb{R}_{>0}$. In other words, we can define a non-linear solution map $G^{D}:\mathcal{F}\to\mathbb{L}^{2}(\mathcal{O})$, mapping any $f\in\mathcal{F}$ to the corresponding solution $u_{f}=G^{D}(f)\in\mathbb{L}^{2}(\mathcal{O})$ of eq. 13. Following eq. 7, we then observe
$Y_{i}=G^{D}(f_{\mathfrak{o}})(X_{i})+\sigma\varepsilon_{i},i\in\left\llbracket N\right\rrbracket,$ (14)
@@ 3 Application: PDE constrained regression models
    $\sup_{x\in\mathbb{R}}\left\{|\Psi^{(k)}(x)|\right\}<\infty,$
@@
For examples and typical choices of (regular) link functions in practice, we refer to Nickl et al. (2020) and Nickl and Wang (2024). In the notation of section 2, throughout this section we set $\mathcal{Z}=\mathcal{O}$, $\mathscr{Z}:=\mathscr{B}_{\mathcal{O}}$ and $\zeta=\mathcal{L}_{\mathcal{O}}$ (recalling that $\operatorname{vol}(\mathcal{O})=1$). We set $\tilde{\Theta}:=\Theta:=\left\{\Psi^{-1}\circ f:\,f\in\mathcal{F}\right\}=H^{\alpha}_{c}(\mathcal{O})$ (the pulled-back) parameter space, and
@@
$J_{\mathfrak{r},N}:\mathcal{F}\to\mathbb{R}_{\leq 0}\cup\{-\infty\},\,f\mapsto-\frac{1}{2\sigma^{2}N}\sum_{i\in\left\llbracket N\right\rrbracket}|Y_{i}-G^{D}(f)(X_{i})|^{2}-\frac{\mathfrak{r}^{2}}{2}\|\Psi^{-1}\circ f\|_{H^{\alpha}(\mathcal{O})}^{2}$ (15)
for any $\mathfrak{r}\in\mathbb{R}_{>0}$. In the notation of eq. 8, we have $J_{\mathfrak{r},N}[f]=\mathscr{J}_{\mathfrak{r},N}[\theta]$ for any $f=f_{\theta}\in\mathcal{F}$, such that maximizing $J_{\mathfrak{r},N}$ over $\mathcal{F}$ is equivalent to maximizing $\mathscr{J}_{\mathfrak{r},N}$ over $\Theta=H^{\alpha}_{c}(\mathcal{O})$. Any pairs of maximizers will be denoted by
@@
$\mathfrak{D}_{\mathfrak{r}}^{2}(f_{1},f_{2}):=\|G(f_{1})-G(f_{2})\|_{\mathbb{L}^{2}_{\zeta}(\mathcal{Z})}^{2}+\mathfrak{r}^{2}\|\Psi^{-1}\circ f_{1}\|_{H^{\alpha}(\mathcal{O})}^{2}.$
@@
Let $\mathcal{F}$ be as in eq. 12 for some $\alpha>1+\frac{5}{2}d$ and $f_{\min}\in\mathbb{R}_{>0}$. Let $G^{D}(f)=u_{f}$ denote the unique solution of eq. 13 and let $D_{N}\sim\mathbb{P}_{f_{\mathfrak{o}}}^{N}$ from eq. 14 for some $f_{\mathfrak{o}}\in\mathcal{F}$. Let $\Psi$ be a regular link function. Let $J_{\mathfrak{r}_{N},N}$ be as in eq. 15 with
@@
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{P}_{f}^{N}\left(\mathfrak{D}_{\mathfrak{r}}^{2}(\hat{f}_{\textsc{MAP}},f)\geq\operatorname{M}\cdot\mathfrak{r}_{N}^{2}\right)\right\}\leq c_{2}\exp\left(-c_{3}\cdot\operatorname{M}^{2}N\mathfrak{r}_{N}^{2}\right).$ (16)
@@
    Assume additionally that $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\bar{c}\in\mathbb{R}_{>0}$, there exists a constant $c_{2}=c_{2}(\alpha,d,\mathcal{O},f_{\min},g_{\min},f_{\mathfrak{o}},\sigma)$ large enough and $c_{3}\in\mathbb{R}_{>0}$, such that
    $\mathbb{P}_{f_{\mathfrak{o}}}^{N}\left(\|\hat{f}_{\textsc{MAP}}-f_{\mathfrak{o}}\|_{\mathbb{L}^{2}(\mathcal{O})}\leq c_{2}\cdot\mathfrak{r}_{N}^{\frac{\alpha-1}{\alpha+1}},\,\|\Psi^{-1}\circ\hat{f}_{\textsc{MAP}}\|_{H^{\alpha}(\mathcal{O})}\leq c_{2}\right)\\ \geq 1-c_{3}\cdot\exp(-c_{1}\cdot N\mathfrak{r}_{N}^{2}).$
@@
Under the conditions formulated in theorem 3.2, we have the following results.
@@
    (Prediction Error) For every $\operatorname{L}\in\mathbb{R}_{>0}$, $\beta\in[0,\alpha+1]$, there exists a finite constant $c_{4}=c_{4}(\alpha,\beta,d,\mathcal{O},f_{\min},L,\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\|G^{D}(\hat{f}_{\textsc{MAP}})-G^{D}(f)\|_{H^{\beta}(\mathcal{O})}^{2}\right]\right\}\leq c_{4}\cdot N^{-\frac{2(\alpha+1-\beta)}{2(\alpha+1)+d}}.$
@@
    (Estimation Error) Let $g\geq g_{\min}\in\mathbb{R}_{>0}$. For every $\operatorname{L}\in\mathbb{R}_{>0}$, there exists a constant $c_{5}=c_{5}(g_{\min},f_{\min},\mathcal{O},d,\alpha,L,\|g\|_{H^{\alpha-1}(\mathcal{O})})\in\mathbb{R}_{>0}$, such that any maximizer $\hat{f}_{\textsc{MAP}}\in\mathcal{F}$ of $J_{\mathfrak{r}_{N},N}$ over $\mathcal{F}$ satisfies
    $\sup_{f\in\mathcal{F}(\operatorname{L})}\left\{\mathbb{E}_{f}^{N}\left[\|\hat{f}_{\textsc{MAP}}-f\|_{\mathbb{L}^{2}(\mathcal{O})}^{2}\right]\right\}\leq c_{5}\cdot N^{-\frac{2(\alpha-1)}{2(\alpha+1)+d}}.$
Remark § 3.4 (Prediction and estimation risk).
  1. i)
    The convergence rates for the prediction risk presented in theorem 3.3 i) align with the classical interpretation of statistical inverse problems. The smoothness of the parameter of interest $f$ is quantified by $\alpha$, while the ill-posedness of the problem is captured by $\kappa=1$, reflecting the smoothing property of the forward map $G^{D}$. Indeed, Lemma 23 in Nickl et al. (2020) shows that if $f$ is $\alpha$-smooth, then $G^{D}(f)$ is $\alpha+1$-smooth. Considering the prediction risk in the $H^{\beta}$-norm, which corresponds to differentiation of order $\beta$, yields the characteristic numerator $\alpha+1-\beta$ in the convergence rate. These rates coincide exactly with those obtained in the white noise model setting in Nickl et al. (2020) for the Darcy problem. Theorem 10 in Nickl et al. (2020) in addition shows that for the $H^{2}$-prediction loss the obtained rate of convergence is minimax-optimal. This optimality also carries over to the random design model considered here, as the proof techniques can be adapted without difficulty. Hence, the results in Nickl et al. (2020) provides a rigorous benchmark indicating that the convergence rates derived in these settings are indeed statistically sharp.
  2. ii)
    The additional assumption $g\geq g_{\min}>0$ in theorem 3.2 iii) and theorem 3.3 ii) plays a crucial role in the analysis. It ensures the validity of a stability estimate for the inverse operator $(G^{D})^{-1}$, which is central to deriving the convergence rate for the estimation error. Such a condition is also imposed in Lemma 24 of Nickl et al. (2020), which serves as a foundation for the current proof strategy. We additionally refer to Richter (1981) and Bonito et al. (2017b) for a broader discussion on injectivity and global stability estimates for inverse problems of this type.
  3. iii)
    In contrast to the prediction error, the rates of convergence for the estimation error derived in theorem 3.3 ii) do not match those obtained in Giordano and Nickl (2020) for the posterior contraction, which are generally slower. This discrepancy is partly due to the lower posterior regularity compared to the MAP estimator, resulting in suboptimal rates for the Bayesian approach. However, the optimality of the estimation error rates for $\hat{f}_{\textsc{MAP}}$ established here remains an open question. Classical nonparametric theory would suggest a faster rate, such as having exponent $\frac{2\alpha}{2(\alpha+1)+d}$. It is unclear to the author whether this gap is a consequence of the estimation methodology and the proof techniques employed—particularly those relying on inverse stability estimates and associated PDE analytic arguments—or whether it reflects an intrinsic structural limitation of the Darcy problem itself. Further investigation is needed to clarify this issue.
The proofs of theorem 3.2 and theorem 3.3 are given in detail in section 4. They are based on verifying that the forward map $\mathscr{G}^{D}$ satisfies C1)-C4) formulated in 2.3 and 2.9 for suitable parameters $\alpha,\kappa,\gamma_{1},\gamma_{2},\eta_{1}$ and $\eta_{2}$, with $V=W=\mathbb{R}$, and an application of the results in section 2.
@@
The general framework presented in section 2 does not only apply to the Darcy problem, it promises to consider various PDE-related problems. As already brielfly discussed in remark 2.4, the results in Bohr and Nickl (2023) concerning the MAP estimate are special cases by setting $\eta_{1}=\eta_{2}=\gamma_{2}=0$. Hence, the solution map of the time-independent Schrödinger equation satisfies the requirements of 2.3 and 2.9 taking the results of Nickl et al. (2020) also into account. Moreover, very analogous and elementary computations show that the solution map of the heat equation studied in Kekkonen (2022) also satisfies the conditions in 2.3 and 2.9 with $\eta_{1}=\eta_{2}=\gamma_{2}=0$.
In very recent work, Nickl and Titi (2024) are considering a non-linear Bayesian data assimilation model for the periodic $d=2$-dimensional Navier Stokes equation with unknown initial condition. In this setting $u=u_{\theta}:\mathcal{O}\times(0,T)\to\mathbb{R}^{2}$ solves the non-linear evolution equation
@@
where $\theta$ is the initial condition of the system, $f$ a forcing term and $\nu\in\mathbb{R}_{>0}$ a viscosity parameter. For this type of equation in two dimensions a complete solution theory already exists. In particular, if the initial condition $\theta$ is contained in some specific function class $\tilde{\mathcal{F}}$, the solution operator $\tilde{\mathcal{F}}\ni\theta\mapsto G^{NS}(\theta)=u_{f}\in\mathbb{L}^{2}(\mathcal{O}\times(0,T))$ is always well-defined. Based on a data set $D_{N}\in(\mathbb{R}^{2}\times\mathcal{O}\times(0,T))^{N}$ arising from eq. 7, it seems feasible to apply the theory from section 2. Indeed, using the theory presented in Nickl and Titi (2024), elementary computations show that the solution map is $(\alpha,0,0)$-regular in $\mathbb{L}^{2}$-norm and uniformly bounded, thus satisfying C1) and C2) from 2.3. C7) from 2.9 can also be shown with $\tau=1$. If C3) of 2.3 is additionally satisfied, one would obtain rates of the form $N^{-\frac{\alpha}{\alpha+2}}$ for estimating $\theta$ - both in prediction and estimation error.
4 Appendix
4.1 Proofs of section 2